OpenStudy (anonymous):

Write an inequality to represent the problem (1 pt.) and then solve the inequality by writing the pairs which solve it (1 pt.). Find all sets of two consecutive positive odd integers whose sum is less than or equal to 18.

OpenStudy (anonymous):

2n + 1) + (2n + 3) <= 18 4n + 4 <=18 4n <= 14 n<= 7/2 Hence, for the smaller of the pair, 2n + 1 <=8, giving the four pairs: (1,3), (3,5), (5,7) and (7,9)